the single displacement reaction K + NaCl ----> Na + KCl

Determine how many grams of sodium will be produced when 20g of K is reacted with an excess of NaCl
1
Expert's answer
2020-08-22T08:25:21-0400

The conversion sequence is;

"gK\to mol K\to molNa\to gNa"

Conversion of mass to moles:

"20gK\times" "1mol K\over 39.0983K" "=0.5115mol K"

Moles to moles conversion( using the mole ratio):

"0.5115mol K\times" "1mol Na\over 1mol K" "=0.5115mol Na"

Moles to mass(g) conversion:

"0.5115mol Na" "\times" "22.9897mol Na\over1mol K" "=11.7592gNa"

The mass of Na produced is "11.7592g"